OpenStudy (anonymous):

What is the midpoint of (3, 4) and (–1, –2)?

OpenStudy (anonymous):

Hello, midpoint of two points \((x_1,y_1)\) and \((x_2,y_2)\) will be\[(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

Use the midpoint formula: (x₁ + x₂ / 2 , y₁ + y₂ / 2) In this case it is: M = [4+(-6) / 2], [(-3+2) / 2] M = (-1, -1/2)

OpenStudy (anonymous):

(1,1)